Volume of CO2 obtained at STP by the complete decomposition of 98.5 g Na2CO3 is

Grade:12th pass

1 Answers

Vikas TU
7 years ago
moles of Na2CO3 = 98.5/106 = 0.9292 mol.
Na2CO3    ---------->         Na2O + CO2
moles of CO2 = 0.9292 mol {From mole 1 analysis}
Volume = 0.9292*22.4 = 20.81 litre.